A city taxi charges $2.50/km. A 6km ride costs $22.
Write the equation that represents the cost, C of a taxi ride compared to the distance driven, d.

Answer by ikleyn(52778)

.

C(d) = (22-6*2.50) + 2.50*d = (22-15) + 2.50*d = 7 + 2.50 *d  dollars.


Here 22-6*50 = 22-15 = 7 dollars is the constant part of the cost.  

     Sometime, they call it "the cost to turn on the counter".


2.50*d is the variable part of the cost.


ANSWER.  The cost function is  C(d) = 7 + 2.50 *d  dollars, where d is the travel distance.

Solved.